A Sample Thesis

With a Subtitle

by Michael McNeil Forbes B.Sc., The University of British Columbia, 1989 M.Sc., The University of Toronto, 1992

A THESIS SUBMITTED IN PARTIAL FULFILMENT OF

THE REQUIREMENTS FOR THE DEGREE OF MASTER OF SCIENCE in The Faculty of Graduate Studies (Department of Physics and Astronomy)

We accept this thesis as conforming to the required standard

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .
. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .
. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .
. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

THE UNIVERSITY OF BRITISH COLUMBIA May 6, 2004 \begin{matrix}c \\ ◯ \end{matrix} Michael McNeil Forbes, 2000

In presenting this thesis in partial fulfilment of the requirements for an advanced degree at the University of British Columbia, I agree that the Library shall make it freely available for reference and study. I further agree that permission for extensive copying of this thesis for scholarly purposes may be granted by the head of my department or by his or her representatives. It is understood that copying or publication of this thesis for financial gain shall not be allowed without my written permission.

(Signature)

Department of Physics and Astronomy

The University Of British Columbia Vancouver, Canada

Date

Abstract

The ubcthesis.cls L^{A}TEX class file and accompanying documents, such as this sample thesis, are distributed in the hope that it will be useful but without any warranty (without even the implied warranty of fitness for a particular purpose). For a description of this file’s purpose, and instructions on its use, see below.

These files are distributed under the GPL which should be included here in the future. Please let the author know of any changes or improvements that should be made.

Michael Forbes. mforbes@physics.ubc.ca

iii

Contents

iv

List of Tables

v

List of Figures Preface

These papers have been published earlier. …

Acknowledgements

Thank you mother here.

Disclaimer

The ubcthesis L^{A}TEX class and the accompanying sample files are unofficial and are not supported by The University of British Columbia. While I have attempted to make the style file and sample files conform to all of the requirements set forth by the Special Collections department of the library, you should always consult one of the Special Collections staff for assistance with problems before starting final draft. You should be able to find the thesis requirements at one of the following sites:

      http://www.library.ubc.ca/
      http://www.library.ubc.ca/spcoll/
      http://www.library.ubc.ca/spcoll/theses.html
      http://www.library.ubc.ca/spcoll/thesesin.html
      http://www.grad.ubc.ca/
      http://www.grad.ubc.ca/currstudents/
      http://www.grad.ubc.ca/currstudents/records/thesis.htm

Table 1: Where to find information regarding thesis preparation at UBC.

A Japanese Introduction

This is a small poem, a little poem, a Haiku, to show you how to. —Michael Forbes.

This small poem shows several features:

• The \newpage command has been used to force a page break.

• The pagestyle has been set to suppress the headers using the command

\thispagestyle{plain}. Note that using \pagestyle{plain} would

have affected all of the subsequent pages.

• The new \chaptertoc{} command has been used to generate a title and to list this “chapter” in the table of contents as “Poem”. If one did not desire to have an entry in the table of contents, then one would just use the starred command \chapter*{}.

• All of this has been done before the \mainmatter command has been issued to delay normal Arabic numbering. Thus, this poem will be numbered with lowercase Roman numerals as part of the front matter.

1

Part I Thesis

Chapter 1

This is a Chapter

1.1 A Section

Here is a section with some text. Equations look like this y = x.

This is an example of a second paragraph in a section so you can see how much it is indented by.

1.1.1 This is a Subsection

Here is an example of a citation: [? ]. The actual form of the citation is governed by the bibliographystyle. These citations are maintained in a BIBTeX file sample.bib. You could type these directly into the file. For an example of the format to use look at the file sample.bbl after you compile this file.

This is an example of a second paragraph in a subsection so you can see how much it is indented by.

This is a Subsubsection

Here are some more citations [? ? ? ]. If you use the natbib package with the sort&compress option, then the following citation will look the same as the first citation in this section: [? ? ? ].

This is an example of a second paragraph in a subsubsection so you can see how much it is indented by.

This is a Paragraph Paragraphs and subparagraphs are the smallest units of text. There is no subsubsubsection etc.

This is a Subparagraph This is the last level of organisation. If you need more than this, you should consider reorganising your work…

f(x) = ∫_{−∞}^{∫_{−∞}^{x} e^{− \frac{y^{2}}{2}} dy} e^{−z^{2}}dz  (1.1)

In order to show you how a separate page would look like (i.e. without a chapter heading) I must type some more text. Thus I will babble a bit and keep babbling for at least one more page… What you should notice is that the chapter titles appear substantially lower than the continuing text. Babble babble babble babble babble babble babble babble babble babble babble babble

Phoenix $960.35
Calgary $250.00

Table 1.1: Here is the caption for this wonderful table.Text of Caption

babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble.

Babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble babble.

Chapter 2

Another Chapter with a Very Long Chapter-name that will Probably Cause Problems

This chapter name is very long and does not display properly in the running headers or in the table of contents. To deal with this, we provide a shorter version of the title as the optional argument to the \chapter[]{} command.

2.1 Another Section

Another bunch of text to demonstrate what this file does. You might want a list for example:

• An item in a list.

• Another item in a list.

An Unnumbered Section That is Not Included in the Table of Contents

Here is an example of a figure environment: Figures can make a document more enjoyable as demonstrated by Figure ??.

An Unnumbered Subsection

Note that if you use subsections or further divisions under an unnumbered section, then you should make them unnumbered as well otherwise you will end up with zeros in the section numbering.

PSfrag replacements π makes me happy!

Figure 2.1: This is a figure of a happy face with a psfrag replacement. The original figure (drawn in xfig and exported to a .eps file) has the text “pie makes me happy!”. The psfrag package replaces this with “π makes me happy!”.

Chapter 3

Landscape Mode

The landscape mode allows you to rotate a page about 90 degrees. Note that this might only work after you convert the dvi file to a postscript (ps) or pdf file using dvips or dvipdf etc.

Appendix A

First Appendix

Here you can have your appendices. Note that the appendix bug will get you here if you use the option upper. Use the modified thesis.cls file to work around this.